Marvel Parts, Inc., manufactures auto accessories. One of the company’s products is a set of seat covers that can be adjusted to fit nearly any small car. The company uses a standard cost system for all of its products. According to the standards that have been set for the seat covers, the factory should work 2,850 hours each month to produce 1,900 sets of covers. The standard costs associated with this level of production are:

Total Per Set
of Covers
Direct materials $ 42,560 $ 22.40
Direct labor $ 51,300 27.00
Variable manufacturing overhead (based on direct labor-hours) $ 6,840 3.60
$ 53.00

During August, the factory worked only 2,800 direct labor-hours and produced 2,000 sets of covers. The following actual costs were recorded during the month:

Total Per Set
of Covers
Direct materials (12,000 yards) $ 45,600 $ 22.80
Direct labor $ 49,000 24.50
Variable manufacturing overhead $ 7,000 3.50
$ 50.80

At standard, each set of covers should require 5.6 yards of material. All of the materials purchased during the month were used in production.

Required:

1. Compute the materials price and quantity variances for August.

2. Compute the labor rate and efficiency variances for August.

3. Compute the variable overhead rate and efficiency variances for August.

(Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Input all amounts as positive values.)
Answer #1

1. Material price variance = (22.40/5.6 - 45600/12000)*12000 = 2400 favorable

Material quantity variance = (2000*5.6-12000)*4 = 3200 unfavorable

2. Labor rate variance = (27*1900/2850 - 49000/2800)*2800 = 1400 favorable

Labor efficiency variance = (2000*2850/1900 - 2800)*18 = 3600 favorable

3. Variable overhead rate variance = (3.6*1900/2850 - 7000/2800)*2800 = 280 unfavorable

Variable overhead efficiency variance = (2000*2850/1900 - 2800)*2.4 = 480 favorable